# 00footnotetext: Received xxxx June xxxx

### v1

@@ 3 Initial data analysis
(a) The $U_{\rm miss}$ distribution of the signal decay $D^{+}\to\gamma e^{+}\nu_{e}$ and (b) the $U^{\prime}_{\rm miss}$ distribution of the validation channel $D^{+}\to\pi^{0}e^{+}\nu_{e}$ after the DT selection. The black dots with error bars represent data. The histograms with filled colors represent different background components in the inclusive MC sample, respectively. The red hollow histogram indicates an arbitrarily scaled signal shape.
liably extract the signal yield from the $U_{\rm miss}$ distribution with adequate sensitivity, thereby requiring further signal identification treatments as elaborated below.
4 Deep-learning-based signal identification
The main challenge in distinguishing the signals from massive backgrounds lies in extracting distinctive features from the data. Notably, the shower patterns observed in the EMC provide key information for this task. Background events involving a $\pi^{0}$ typically result in a higher number of photon candidates than the signal, whereas $K^{0}_{L}$-induced showers often exhibit more dispersed cluster shapes compared to those from photons [42]. Additionally, the kinematics of the particles, influenced by the unique phase spaces and dynamics of their decays, offer another aspect of differentiation. Conventional cut-based methods characterize such distinctions via hand-crafted variables derived from practical understanding of the involved physics, such as the number of $\pi^{0}$-like photon combinations, the lateral moment of photon candidates [20], and the $\chi^{2}$ values from kinematic fits under specific final-state hypotheses [43]. In contrast, deep learning methods aim to process minimally refined data at the level of fundamental detector responses to the penetrations of the final state particles. The DNN is expected to automatically explore underlying correlations and recognize distinctive information beyond prior knowledge.
The dataset for training the DNN is constructed by randomly shuffling signal and background events sourced from the relevant MC samples. The events are categorized into three groups with equal statistics: the $D^{+}\to\gamma e^{+}\nu_{e}$ signal, the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background, and other inclusive backgrounds. The total training dataset contains approximate 3.3 million events, with 80% used for training and the remaining 20% reserved for validation. The input information from an event to DNN includes all charged tracks reconstructed in the MDC and isolated showers clustered in the EMC. These particles are organized as a point cloud [44] structure, i.e., an unordered and variable-sized set of the points in a high-dimensional feature space. For each charged track, the features comprise the azimuthal and polar angles in the center-of-mass frame, the charge, the magnitude of momentum, and the parameters characterizing its helical trajectory in the MDC. Additionally, low-level measurements from the MDC, TOF, and EMC are incorporated as implicit information for particle identification and reconstruction quality. For each shower, the features consist of the azimuthal and polar angles, the energy deposition, the count of fired crystals, the time measurement, and the parameters describing its expansion scope among nearby crystals. A full list of input features is provided in Table 4, and these features are found to be consistent between data and MC simulation.
[…]
@@ 4 Deep-learning-based signal identification
The input particle feature variables and their availability in tracks and showers. If any measurement is unavailable from the detector readouts, the default value is set at zero. Variable Definition Track Shower $\theta$ azimuthal angle $\phi$ polar angle $p$ momentum magnitude charge electric charge $d_{\rho}$ distance in the $x-y$ plane from the point of closest approach (POCA) of the track to the IP $d_{Z}$ distance along $z$-axis from the POCA of the track to the IP $\phi_{0}$ azimuthal angle of the track at the POCA $\kappa$ scaled inverse of the track’s momentum $\tan\lambda$ tangent value of the track’s angle at the POCA with respect to the $x-y$ plane $\chi_{{\rm d}E/{\rm d}x}(e)$ deviation in ${\rm d}E/{\rm d}x$ measurement of the track from the anticipated positron value $\chi_{{\rm d}E/{\rm d}x}(\pi)$ deviation in ${\rm d}E/{\rm d}x$ measurement of the track from the anticipated pion value $\chi_{{\rm d}E/{\rm d}x}(K)$ deviation in ${\rm d}E/{\rm d}x$ measurement of the track from the anticipated kaon value $\chi_{\rm TOF}(\pi)$ deviation in TOF measurement of the track from the anticipated pion value $\chi_{\rm TOF}(K)$ deviation in TOF measurement of the track from the anticipated kaon value $Q_{\rm TOF}$ pulse height of the track in TOF corresponding to its deposited energy $\beta_{\rm TOF}$ velocity of the track measured by TOF $E_{\rm raw}$ total deposited energy in EMC $E_{\rm seed}$ deposited energy of the crystal at the shower center $E_{3\times 3}$ sum of energies of $3\times 3$ crystal matrices surrounding the shower center $E_{5\times 5}$ sum of energies of $5\times 5$ crystal matrices surrounding the shower center $N_{\rm hits}$ count of fired crystals in the shower time EMC time measurement secondary moment shower expansion parameter defined as $\sum_{i}E_{i}r^{2}_{i}/\sum_{i}E_{i}$ lateral moment shower expansion parameter defined as $\sum_{i=3}^{n}E_{i}r_{i}^{2}/(E_{1}r_{0}^{2}+E_{2}r_{0}^{2}+\sum_{i=3}^{n}E_{i}r_{i}^{2})$ $A_{20}$ moment shower expansion parameter defined as $\sum_{i}\frac{E_{i}}{E_{tot}}f_{2,0}(\frac{r_{i}}{R_{0}})$, $f_{2,0}(x)=2x^{2}-1$ $A_{42}$ moment same but defined as $\left|\sum_{i}\frac{E_{i}}{E_{tot}}f_{4,2}(\frac{r_{i}}{R_{0}})e^{2i\phi}\right|$$f_{4,2}(x)=4x^{4}-3x^{2}$
@@
The architecture of the DNN largely inherits from the Particle Transformer (ParT) [45], with adaptations tailored for the BESIII experiment. A key innovation of the ParT is the injection of particle physics knowledge into the Transformer network [46], where a set of pairwise interaction features are calculated using the energy-momentum vectors for each pair of particles and incorporated as biases $\mathbf{U}$ in the self-attention mechanism. This treatment efficiently captures the kinematic relationships between particles, significantly improving the ParT performance. For the BESIII application, we account for its unique properties, such as the purely spherical coordinate system and energy-momentum conservation in 3D space, by designing a new set of pairwise interaction features $\mathbf{U}_{\rm BES}$ as
@@
where $E_{a}$ and $E_{b}$ are the energies of the two particles, $\vec{p}_{a}$ and $\vec{p}_{b}$ are their momentum vectors, and $\theta_{a,b}$ is the opening angle between these vectors in the center-of-mass frame. For the model hyperparameters, this adaptation includes four particle attention blocks and two class attention blocks. It uses a particle embedding encoded from the input particle features via a 3-layer Multilayer Perceptron with 128, 256, 128128256128128,\ 256,\ 128128 , 256 , 128 nodes in each layer, and an interaction embedding encoded from the input pairwise features using a 4-layer pointwise 1D convolution with 64, 64, 64, 8646464864,\ 64,\ 64,\ 864 , 64 , 64 , 8 channels. A dropout rate of 0.1 is applied to prevent overfitting.
The training of the DNN is performed on the aforementioned dataset for 50 epochs with a batch size of 512 and an initial learning rate of 0.001. An additional treatment is employed during training to address the potential correlation between the DNN output and the $U_{\rm miss}$ distribution, which could inadvertently create signal-like structures in the $U_{\rm miss}$ distribution of remaining background events after applying the selection criteria on the DNN output [47, 48, 49, 50]. To decouple the DNN output from the $U_{\rm miss}$ observable, we employ an iterative approach by assigning a weight $\omega(U_{\rm miss})$ to each background event which adjusts its contribution to the loss function. Events with higher weights have a greater impact on the loss function, thus being better classified during training. The weight $\omega_{i}^{j}(U_{\rm miss}^{j})$ for the event $j$ in the $i$-th iteration is calculated as
@@
where $\omega_{0}^{j}(U_{\rm miss}^{j})=1$, $p^{\rm BKG}_{i-1}(U_{\rm miss})$ represents the normalized probability density function for the remaining background shape after the $(i-1)$-th iteration, and $p^{\rm BKG}_{\rm orig}(U_{\rm miss})$ denotes the original background shape before applying the DNN output veto. For each iteration, the DNN is re-trained, and both $p^{\rm BKG}_{i-1}(U_{\rm miss})$ and $\omega_{i}^{j}(U_{\rm miss}^{j})$ are updated. Intuitively, the process converges when $p^{\rm BKG}_{i-1}(U_{\rm miss})$ approaches $p^{\rm BKG}_{\rm orig}(U_{\rm miss})$, resulting in a trivial background shape that facilitates signal extraction. In practice as illustrated in Figure 4, this iterative approach partially corrects the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background shape, while successfully retrieves the original for other inclusive background shape. This result demonstrates both its effectiveness and limitations in addressing the mass correlation issue. Furthermore, another machine-learning technique called model ensemble is employed. With some randomness factors incorporated in the training such as network initialization, batch processing sequence and dropout [51], we train a total of 20 DNNs in parallel in each iteration. The outputs of these DNNs are averaged for each event during inference, creating an ensemble DNN that offers better robustness and generalization than any single DNN model.
@@
The DNN output assigns three scores between [0, 1]01[0,\,1][ 0 , 1 ] to each event, reflecting the probabilities of the event belonging to the signal, $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background and other background categories. Since the sum of the three scores always equals one, only two of them are independent. As the final step in the event selection, we require the score for the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background to be less than 0.15 and the score for other backgrounds to be less than 0.05. These cut values are optimized using the Punzi significance metric, $\epsilon/(1.5+\sqrt{B})$ [52], where $\epsilon$ represents the signal efficiency and $B$ is the expected background yield. Figure 4(a) illustrates the $U_{\rm miss}$ distribution after implementing the deep learning approach. The backgrounds are reduced by more than two orders of magnitude at the cost of approximately two-thirds of the signal efficiency, significantly enhancing the sensitivity of the signal search.
@@
(a) The $U_{\rm miss}$ distribution of the signal decay $D^{+}\to\gamma e^{+}\nu_{e}$ and (b) the $U^{\prime}_{\rm miss}$ distribution of the validation decay $D^{+}\to\pi^{0}e^{+}\nu_{e}$ after the DNN vetoes. The black dots with error bars represent data. The histograms with filled colors represent different background components in the inclusive MC sample, respectively. The red hollow histogram indicates an arbitrarily scaled signal shape.
@@ 5 Signal extraction
According to Eq. (2), the DT efficiencies for each tag mode, $\epsilon^{i}_{\rm DT}$, are estimated using signal MC sample and are summarized in Table 3. To obtain the DT yield $N_{\rm DT}$ in data, a maximum likelihood fit is performed on the $U_{\rm miss}$ distribution combining all the tag modes. The fit includes three components, namely, the signal, the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background and other backgrounds. The shapes for these components are taken from the corresponding MC simulations, and their yields are all floated in the fit. The fit gives $N_{\rm DT}=12.2\pm 8.6$ and $\mathcal{B}(D^{+}\to\gamma e^{+}\nu_{e})=(0.54\pm 0.38)\times 10^{-5}$ as shown in Figure 5(a), where the uncertainties are statistical only. Since no significant signal is observed, an upper limit on the signal BF will be set after accounting for all systematic uncertainties.
@@ 6 Validation
A common concern in adopting the deep learning method in particle physics experiments is the coupling between potential analysis biases and the sophisticated DNN structure, which makes these biases difficult to validate, quantify and calibrate. The extensive and diverse datasets collected at BESIII facilitate various control samples in data to investigate these issues. In this study, our primary background process $D^{+}\to\pi^{0}e^{+}\nu_{e}$ serves as a natural control sample. By partially reversing the DNN vetoes, i.e., to require the score for the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background to be greater than 0.15 while keeping the score for other backgrounds below 0.05, we can isolate a $D^{+}\to\pi^{0}e^{+}\nu_{e}$ sample with a purity of over 98% and measure its absolute BF. Figures 3(b) and 4(b) illustrate the $U^{\prime}_{\rm miss}$ distributions before and after applying the partially reversed DNN vetoes, with significant subtraction of backgrounds other than $D^{+}\to\pi^{0}e^{+}\nu_{e}$.
To measure the absolute BF of $D^{+}\to\pi^{0}e^{+}\nu_{e}$, we use the $\epsilon^{i}_{\rm DT}$ estimated from MC simulations and extract the $N_{\rm DT}$ via a maximum likelihood fit shown in Figure 5(b). In the fit, the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ signal is modeled using the MC simulated shape convolved with a Gaussian resolution function. The other background component is described with the inclusive MC shape with the yield floated. The final fit to all data gives $N_{\rm DT}=3004\pm 59$ and $\mathcal{B}(D^{+}\to\pi^{0}e^{+}\nu_{e})=(3.629\leavevmode\nobreak\ \pm 0.071)\leavevmode\nobreak\ \times 10^{-3}$ where the uncertainty is statistical only. This result is consistent with the previous BESIII measurement $(3.63\leavevmode\nobreak\ \pm 0.08_{\rm stat.}\leavevmode\nobreak\ \pm 0.05_{\rm syst.})\times 10^{-3}$ [53] based on about one sevenths of full data, thereby validating the reliability of our DNN method. The residual deviations are considered as a source of systematic uncertainty, as detailed in the next section.
@@ 7 Systematic uncertainties
The systematic uncertainties in the BF determination stem from various sources, with many related to the ST selection being cancelled benefiting from the DT technique. The remaining systematic sources can be grouped into two categories: multiplicative uncertainties, which affect the efficiency, and additive uncertainties, which impact the signal yield. These two types of uncertainties contribute to the upper limit of the BF through different approaches.
The multiplicative uncertainty sources include the ST yield, positron tracking, positron PID, photon reconstruction, MC model, and DNN vetoes, as summarized in Table 7. The uncertainty due to ST yield is estimated by varying the descriptions of signal and background components in the $M_{\rm BC}$ fits and is assigned as 0.3%. A control sample of radiative Bhabha scattering events is used to study the position tracking and PID uncertainties, and their values are assigned as 1.0% for both [54]. The photon reconstruction uncertainty is assigned to be 1.0% using a control sample of $J/\psi\to\pi^{+}\pi^{-}\pi^{0},\,\pi^{0}\to\gamma\gamma$ [55]. The uncertainty related to MC modeling is assigned as the difference in the signal efficiencies of the nominal MC model with an alternative single-pole model [18], which is 9.1%.
Special considerations are given to systematic uncertainties related to the DNN vetoes [56]. As suggested by some frontier surveys [13, 57], we consider two primary uncertainty sources: model uncertainty and domain shift. Model uncertainty arises from our limited knowledge of the best model and is quantified by the relative shift in signal efficiency across the 20 DNNs trained during the model ensemble process. The standard deviation of these efficiencies, divided by their average, 5.3%, is taken as the corresponding systematic uncertainty. Domain shift describes the mismatch between datasets for training and inference, reflecting potential data-MC inconsistencies. This uncertainty is evaluated using the control sample of $D^{+}\to\pi^{0}e^{+}\nu_{e}$ described in Section 6, with the deviation between our BF measurement and the previous BESIII measurement [53] calculated to be 2.2%, combining the center value shift and $1\sigma$ statistical uncertainty. In total, the multiplicative systematic uncertainty is determined to be 10.9% by adding the above sources in quadrature.
@@
The additive systematic uncertainties arise from the signal, $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background and other background shapes used in the DT fit. These effects are evaluated by testing alternative descriptions of the shapes. One variation is using parameterized functions, where the signal and $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background shapes are modeled by double-Gaussian functions, and the other background shape by a 3rd-order Chebyshev polynomial. Another change is using the MC-simulated shapes before the final iteration of the mass decorrelation process, as well as across the 20 DNN predictions in the model ensemble process.

### v2

@@ 4 Deep-learning-based signal identification
Input particle feature variables and their availability in tracks and showers. If measurements were unavailable from the detector readouts, the default value was set to zero. Variable Definition Track Shower $\theta$ azimuthal angle $\phi$ polar angle $p$ momentum magnitude charge electric charge $d_{\rho}$ distance in the xyx-y plane from the point of closest approach (POCA) of the track to the IP $d_{Z}$ distance along $z$-axis from the POCA of the track to the IP $\phi_{0}$ azimuthal angle of the track at the POCA $\kappa$ scaled inverse of the momentum of the track $\tan\lambda$ tangent value of the angle of the track at the POCA with respect to the xyx-y plane $\chi_{{\rm d}E/{\rm d}x}(e)$ deviation in ${\rm d}E/{\rm d}x$ measurement of the track from the anticipated positron value $\chi_{{\rm d}E/{\rm d}x}(\pi)$ deviation in ${\rm d}E/{\rm d}x$ measurement of the track from the anticipated pion value $\chi_{{\rm d}E/{\rm d}x}(K)$ deviation in ${\rm d}E/{\rm d}x$ measurement of the track from the anticipated kaon value $\chi_{\rm TOF}(\pi)$ deviation in TOF measurement of the track from the anticipated pion value $\chi_{\rm TOF}(K)$ deviation in TOF measurement of the track from the anticipated kaon value $Q_{\rm TOF}$ pulse height of the track in TOF corresponding to its deposited energy $\beta_{\rm TOF}$ velocity of the track measured by TOF $E_{\rm raw}$ total deposited energy in EMC $E_{\rm seed}$ deposited energy of the crystal at the shower center $E_{3\times 3}$ sum of energies of $3\times 3$ crystal matrices surrounding the shower center $E_{5\times 5}$ sum of energies of $5\times 5$ crystal matrices surrounding the shower center $N_{\rm hits}$ count of fired crystals in the shower time EMC time measurement secondary moment shower expansion parameter defined as $\sum_{i}E_{i}r^{2}_{i}/\sum_{i}E_{i}$ lateral moment shower expansion parameter defined as $\sum_{i=3}^{n}E_{i}r_{i}^{2}/(E_{1}r_{0}^{2}+E_{2}r_{0}^{2}+\sum_{i=3}^{n}E_{i}r_{i}^{2})$ $A_{20}$ moment shower expansion parameter defined as $\sum_{i}\frac{E_{i}}{E_{tot}}f_{2,0}(\frac{r_{i}}{R_{0}})$, $f_{2,0}(x)=2x^{2}-1$ $A_{42}$ moment same but defined as $\left|\sum_{i}\frac{E_{i}}{E_{tot}}f_{4,2}(\frac{r_{i}}{R_{0}})e^{2i\phi}\right|$, $f_{4,2}(x)=4x^{4}-3x^{2}$
@@
The architecture of the DNN used in this study largely inherits from the Particle Transformer (ParT) [45], with adaptations tailored for the BESIII experiment. A key innovation of the ParT is the injection of particle physics knowledge into the Transformer network [46], where a set of pairwise interaction features are calculated using the energy-momentum vectors for each pair of particles and incorporated as biases $\mathbf{U}$ in the self-attention mechanism. This treatment efficiently captures the kinematic relationships between particles, significantly improving the ParT performance. For the BESIII application, we considered its unique properties, such as the purely spherical coordinate system and energy-momentum conservation in 3D space, by designing a new set of pairwise interaction features $\mathbf{U}_{\rm BES}$ defined as
@@
where $E_{a}$ and $E_{b}$ are the energies of the two particles, $\vec{p}_{a}$ and $\vec{p}_{b}$ are their momentum vectors, and $\theta_{a,b}$ is the opening angle between these vectors in the center-of-mass frame. For the model hyperparameters, this adaptation includes four particle attention blocks and two class attention blocks. It uses a particle embedding encoded from the input particle features via a 3-layer multilayer perceptron with 128, 256, 128128,\ 256,\ 128 nodes in each layer and an interaction embedding encoded from the input pairwise features using a 4-layer pointwise 1D convolution with 64, 64, 64, 864,\ 64,\ 64,\ 8 channels. A dropout rate of 0.1 was applied to prevent overfitting.
The training of the DNN was performed on the aforementioned dataset for 50 epochs with a batch size of 512 and an initial learning rate of 0.001. An additional treatment was employed during training to address the potential correlation between the DNN output and the $U_{\rm miss}$ distribution, which could inadvertently create signal-like structures in the $U_{\rm miss}$ distribution of remaining background events after applying the selection criteria on the DNN output [47, 48, 49, 50]. To decouple the DNN output from the $U_{\rm miss}$ observable, we employed an iterative approach by assigning a weight $\omega(U_{\rm miss})$ to each background event which adjusts its contribution to the loss function. Events with higher weights have a greater impact on the loss function, which means that they are better classified during training. The weight $\omega_{i}^{j}(U_{\rm miss}^{j})$ for an event $j$ in the $i$-th iteration is calculated as
@@
where $\omega_{0}^{j}(U_{\rm miss}^{j})=1$, $p^{\rm BKG}_{i-1}(U_{\rm miss})$ represents the normalized probability density function for the remaining background shape after the (i1)(i-1)-th iteration, and $p^{\rm BKG}_{\rm orig}(U_{\rm miss})$ denotes the original background shape before applying the DNN output veto. For each iteration, the DNN was re-trained, and both $p^{\rm BKG}_{i-1}(U_{\rm miss})$ and $\omega_{i}^{j}(U_{\rm miss}^{j})$ were updated. Intuitively, the process converged when $p^{\rm BKG}_{i-1}(U_{\rm miss})$ approached $p^{\rm BKG}_{\rm orig}(U_{\rm miss})$, resulting in a trivial background shape that facilitates signal extraction. In practice, as shown in Figure 4, this iterative approach partially corrects the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background shape while successfully retrieves the original one for other inclusive background shape. This result demonstrates both its effectiveness and limitations in addressing the mass correlation issue. Furthermore, another machine-learning technique called model ensemble was employed. With some randomness factors incorporated in the training, such as network initialization, batch processing sequence, and dropout [51], we trained a total of 20 DNNs in parallel in each iteration. The outputs of these DNNs were averaged for each event during inference, creating an ensemble DNN that offers better robustness and generalization than any single DNN model.
@@
The DNN output assigns three scores in the range [0, 1][0,\,1] to each event, reflecting the probabilities of the event belonging to the signal, $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background and other background categories. Given that the sum of the three scores always equals one, only two of them are independent. As the final step in the event selection, we required the score for the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background to be less than 0.15 and that for other backgrounds to be less than 0.05. These cut values were optimized using the Punzi significance metric, $\epsilon/(1.5+\sqrt{B})$ [52], where $\epsilon$ represents the signal efficiency and $B$ is the expected background yield. Figure 4(a) shows the $U_{\rm miss}$ distribution after implementing the deep learning approach. The backgrounds were reduced by more than two orders of magnitude at the cost of approximately two-thirds of the signal efficiency, significantly enhancing the sensitivity of the signal search.
@@
(a) $U_{\rm miss}$ distribution of the signal decay $D^{+}\to\gamma e^{+}\nu_{e}$ and (b) $U^{\prime}_{\rm miss}$ distribution of the validation decay $D^{+}\to\pi^{0}e^{+}\nu_{e}$ after the DNN vetoes. The black dots with error bars represent data. The histograms with filled colors represent different background components in the inclusive MC sample. The red hollow histogram indicates an arbitrarily scaled signal shape.
@@ 5 Signal extraction
According to Eq. (2), the DT efficiencies for each tag mode, $\epsilon^{i}_{\rm DT}$, were estimated using signal MC sample; they are summarized in Table 3. To obtain the DT yield $N_{\rm DT}$ in data, a maximum likelihood fit was performed on the $U_{\rm miss}$ distribution combining all the tag modes. The fit included three components, namely the signal, $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background, and other backgrounds. The shapes for these components were taken from the corresponding MC simulations, and their yields were all floated in the fit. The fit resulted in $N_{\rm DT}=12.2\pm 8.6$ and $\mathcal{B}(D^{+}\to\gamma e^{+}\nu_{e})=(0.54\pm 0.38)\times 10^{-5}$, as shown in Figure 5(a), where the uncertainties are statistical only. Given that no significant signal was observed, an upper limit on the signal BF was set after accounting for all systematic uncertainties.
@@ 6 Validation
A common concern in adopting the deep learning method in particle physics experiments is the coupling between potential analysis biases and the sophisticated DNN structure, which makes these biases difficult to validate, quantify, and calibrate. The extensive and diverse datasets collected at BESIII facilitate various control samples in data to investigate these issues. In this study, our primary background process $D^{+}\to\pi^{0}e^{+}\nu_{e}$ serves as a natural control sample. By partially reversing the DNN vetoes, i.e., to require the score for the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background to be greater than 0.15 while keeping the score for other backgrounds below 0.05, we can isolate a $D^{+}\to\pi^{0}e^{+}\nu_{e}$ sample with a purity exceeding 98% and measure its absolute BF. Figures 3(b) and 4(b) show the $U^{\prime}_{\rm miss}$ distributions before and after applying the partially reversed DNN vetoes, with significant subtraction of backgrounds other than $D^{+}\to\pi^{0}e^{+}\nu_{e}$.
To measure the absolute BF of $D^{+}\to\pi^{0}e^{+}\nu_{e}$, we used the value of $\epsilon^{i}_{\rm DT}$ estimated from MC simulations and extract $N_{\rm DT}$ via the maximum likelihood fit shown in Figure 5(b). In the fit, the $D^{+}\to\pi^{0}e^{+}\nu_{e}$ signal was modeled using the MC simulated shape convolved with a Gaussian resolution function. The other background component was described using the inclusive MC shape with the yield floated. The final fit to all data resulted in $N_{\rm DT}=3004\pm 59$ and $\mathcal{B}(D^{+}\to\pi^{0}e^{+}\nu_{e})=(3.629\penalty 10000\ \pm 0.071)\penalty 10000\ \times 10^{-3}$, where the uncertainty was statistical only. This result is consistent with previous BESIII measurement $(3.63\penalty 10000\ \pm 0.08_{\rm stat.}\penalty 10000\ \pm 0.05_{\rm syst.})\times 10^{-3}$ [53], based on about one sevenths of full data, thereby validating the reliability of our DNN method. The residual deviations were considered a source of systematic uncertainty, as detailed in the next section.
@@ 7 Systematic uncertainties
Systematic uncertainties in BF determination stem from various sources; many of them are related to the ST selection being canceled, benefiting from the DT technique. The remaining systematic sources can be grouped into two categories: multiplicative uncertainties, which affect the efficiency, and additive uncertainties, which impact the signal yield. These two types of uncertainties contribute to the upper limit of the BF through different approaches.
The multiplicative uncertainty sources include the ST yield, positron tracking, positron PID, photon reconstruction, MC model, and DNN vetoes, as summarized in Table 7. The uncertainty due to the ST yield was estimated by varying the descriptions of signal and background components in the $M_{\rm BC}$ fits and was set to 0.3%. A control sample of radiative Bhabha scattering events was used to study the position tracking and PID uncertainties, and the values for both were set to 1.0% [54]. The photon reconstruction uncertainty was set to 1.0% using a control sample of $J/\psi\to\pi^{+}\pi^{-}\pi^{0},\,\pi^{0}\to\gamma\gamma$ [55]. The uncertainty related to MC modeling was defined as the difference between the signal efficiencies of the nominal MC model and an alternative single-pole model [18]; it was set to 9.1%.
Special consideration was given to systematic uncertainties related to the DNN vetoes [56]. Following some frontier surveys [13, 57], we consider two primary uncertainty sources: model uncertainty and domain shift. Model uncertainty arises from our limited knowledge of the best model and was quantified by the relative shift in signal efficiency across the 20 DNNs trained during the model ensemble process. The standard deviation of these efficiencies, divided by their average, 5.3%, was taken as the corresponding systematic uncertainty. Domain shift describes the mismatch between datasets for training and inference, reflecting potential data-MC inconsistencies. This uncertainty was evaluated using the control sample of $D^{+}\to\pi^{0}e^{+}\nu_{e}$ described in Section 6, with the deviation between our BF measurement and the previous BESIII measurement [53] calculated to be 2.2%, combining the center value shift and $1\sigma$ statistical uncertainty. In total, the multiplicative systematic uncertainty was determined to be 10.9% by adding the above sources in quadrature.
@@
The additive systematic uncertainties arise from the signal, $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background, and other background shapes used in the DT fit. These effects were evaluated by testing alternative descriptions of the shapes. One variation is in using parameterized functions, where the signal and $D^{+}\to\pi^{0}e^{+}\nu_{e}$ background shapes were modeled by double-Gaussian functions, and the other background shape was modeled a third-order Chebyshev polynomial. Another change was the use of MC-simulated shapes before the final iteration of the mass decorrelation process as well as across the 20 DNN predictions in the model ensemble process.